e.g.23 An electric motor of 10 kg is suffering from vibration due to unbalance of 0.01 kg. m2. The stiffness coefficient of tclear handwriting , no typing , show all work please!
Answer #1

con Tos 낙 100 rad/s 10 when. we take. 2 = 0.1 wna w. J-22 loo - 100.60 radis - z² 1-(0.172 4 Now, c = 2zon = 20 = 2(0.1) 1001 oo 19 Now, 4 when • z = 0.5 = 116.27 rodis JI-(0.5) 2 = 2 22n = 2(0.5) 100 = loo fi = 5K² + c²w2 J GOS) + (coo)? x 16.27) ?3 so for the minimum transmitted force (F) we take damping factor ? = 0.1 ans.